OpenStudy (anonymous):

Solve the given linear system y=5x-3 y=3x-1

OpenStudy (anonymous):

first equation minus 2nd equation gives 2x-2 =0 x =1 y = 3*3-1 =2

OpenStudy (anonymous):

5x-3=3x-1 2x=-1+3 2x=2 x=1

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

w

OpenStudy (anonymous):

e

OpenStudy (anonymous):

l

OpenStudy (anonymous):

c

OpenStudy (anonymous):

o

OpenStudy (anonymous):

m

OpenStudy (anonymous):

e

OpenStudy (anonymous):

combo breaker

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

can i ask another one?

OpenStudy (anonymous):

-2x+3y=8 3x+5y=-12?

OpenStudy (anonymous):

http://www.hellam.net/algebra/simul1.html

OpenStudy (anonymous):

no solution :(

OpenStudy (anonymous):

http://www.webmath.com/solver2.html

OpenStudy (anonymous):

xD thanks sir.

OpenStudy (anonymous):

Hey i m 13 by the way

OpenStudy (anonymous):

Not a sir lol